# Majority Voting of Doctors Improves Appropriateness of AI Reliance in Pathology

### v1

@@ 4.5.1 Utilization of AI & XAI (RQ1)
    XAI activation rate was calculated according to Equation 3. The number of “AI recommendations in ${}_{i}$” was counted based on the highest sensitivity set by a participant while they examined the ${}_{i}$. If the “Show AI” was not toggled on in an image, then it was not counted.
@@ 4.5.2 Reliance on AI (RQ2)
Items
Majority Voting Decisions (Group Size=$k$)
Human Decision (stage 1)
Majority voting results based on the stage 1 decisions from $k$ participants
AI Recommendation (stage 2)
For each image, AI recommendations under the highest sensitivity set by more than $k/2$ of participants while they were seeing the image
Human-AI Decision (stage 2)
Majority voting results based on the stage 2 decisions from the same $k$ participants
@@ 4.5.3 Correctness of Mitosis Detection (RQ3)
Here, we compare the precision and recall of five conditions: one-human-AI collaboration, and majority voting decisions from AI-assisted pathologists (group sizes $k$=3,5,7,9). Results of larger group sizes are reported in B. Similar to the comparisons in the AI reliance metrics, for each of precision and recall, a Kruskal–Wallis and a follow-up post-hoc Dunn’s test with Bonferroni correction was employed to test the significance among the condition pairs.
Because our previous work showed AI alone outperformed all participants in stage 1 (Gu et al., 2024), the “complementary team performance” in this work refers explicitly to cases where the human+AI approach outperforms AI (RQ3). Here, the AI operating point was selected based on the best threshold in the model validation process. For precision and recall, we defined the “success rate of achieving complementary team performance” using Equation 9. This equation was applied to both the one-human-AI collaboration, and majority voting decision conditions with group sizes $k$ ranging from 3 to 27.
@@ 5.1 Utilization of AI & XAI
The mean AI activation rate was $M=99.21\%$ ($SD=0.481\%$, $CI_{95}=[98.13\%,100.00\%]$, Figure 6(a)) [^6] The mean ($M$), standard deviation ($SD$), and 95% confidence intervals ($CI_{95}$) were calculated by the bootstrapping method (100% re-sampling with replacement, 10,000 times) . And the mean AI active time percentage was $M=71.39\%$ ($SD=6.713\%$, $CI_{95}=[57.75\%,83.94\%]$, Figure 6(b)). 21/29 participants had $>50\%$ AI active time percentages, with an example of how they interacted with the “Show AI” feature shown in Figure 6(c). Interestingly, the remaining 8/29 participants had $<25\%$ AI active time percentages: although the “Show AI” feature was majority deactivated, these participants would still activate AI recommendations briefly while examining each image (Figure 6(d)).
@@ 5.3 Correctness of Mitosis Detection
Figure 8: Box-whisker plots of precision and recall for the five conditions of one-human-AI collaboration, and majority voting decisions ($k=3,5,7,9$); (c) Precision-recall plots for mitosis detection for these five conditions. The red line represents the precision-recall curve of AI, and the ‘x’ marker indicates the AI’s performance at a threshold determined by the best validation performance. The success rates of achieving complementary team performance (i.e., percentage of human+AI cases where their performance is higher than both humans and AI) for the criteria of (d) precision and (e) recall.
@@
Regarding the success rates for achieving complementary team performance, for precision, none of the majority voting conditions (i.e., $k=3\rightarrow 27$) was higher than the success rate achieved by one-human-AI collaboration ($13.87\%$ success rate, Figure 8(d)). On the other hand, for recall, all majority voting conditions had higher success rates compared to one-human-AI collaboration ($51.72\%$ success rate): As shown in Figure 8(e), the lowest success rate was observed at $k=3$ ($53\%$ success rate), and the highest was achieved at $k=27$, reaching $76\%$.
@@ 6.1.1 Summary of RQ1
AI was activated at least once in most images. However, this does not imply that the AI was constantly active throughout the entire study. Notably, 8/29 participants deactivated AI for most of the study, and only activated it briefly occasionally. That is, in certain instances, the ‘AI on-request’ feature posed cognitive forcing function effects.
@@ 6.1.3 Summary of RQ3
All majority voting conditions for $k=3\rightarrow 27$ did not show a higher success rate in achieving complementary precision than one-human-AI collaboration. This is because AI had a high precision of 0.961, and there was a lack of space for improvement. For recall, all majority voting conditions ($k=3\rightarrow 27$) showed higher success rates. Notably, the highest success rate, 76%, was achieved at $k=27$, indicating a 46.95% increase over the one-human-AI collaboration condition (51.72% success rate).
Figure 9: Bar plots for the agreements among 29 participants for (a) 88 ground-truth mitoses, and (b) 91 false-positive mitoses that at least two participants agreed on. The diamond markers ($\lozenge$) stand for the AI detections under the “Highest” AI sensitivity setting. 1⃝ An example of under-reliance that might not be addressed by the majority voting; 2⃝ An example of under-reliance that might be addressed by the majority voting; 3⃝ An example of over-reliance that might not be addressed by the majority voting; and 4⃝ An example of over-reliance that might be addressed by the majority voting.
@@ 6.2 The Mechanism and Cost of Majority Voting
    Reducing Over-Reliance on AI False Positives: AI’s false-positive detections led to higher agreement rates among participants (as shown in Figure 9(b)3⃝), suggesting participants’ tendency of over-reliance in at stage 2. The majority of these false-positive detections did not achieve agreement rates higher than 50% (Figure 9(b)4⃝). In other words, from a group’s perspective, it was not usual for the majority of participants to consistently over-rely when AI made false-positive mistakes. Therefore, the over-reliance can be reduced by the majority voting mechanism.
@@
    Reducing Under-Reliance on Human False Positives: At stage 2, participants may make the same false-positive mistake as in stage 1, even when AI correctly suggested negative (Figure 9(b)1⃝). This suggests that the under-reliance incidents happened when one participant collaborated with AI. Nevertheless, agreement rates for these false positives rarely exceeded the 50% majority threshold (Figure 9(b)2⃝), indicating that majority voting could reduce under-reliance.
@@ 6.4 Towards Efficient & Reliable Medical Decisions with AI+k
Determining the optimal balance between decision-making and time expenditure has been well-explored in previous crowd-sourcing works (Daniel et al., 2018). However, one should be aware that the workflow of medical professionals is usually different from that of general users, and their preferences in using AI and XAI may also vary (as shown in Section 5.1). Therefore, future research should focus on exploring which AI+$k$ methods can seamlessly integrate into the workflow of medical professionals, effectively balancing efficiency and reliability in the medical decisions of multiple doctors.
@@ 7 Conclusion
This study introduces and validates the majority voting approach to enable doctors’ appropriate reliance on medical AI. By recruiting 32 pathology professionals, we conducted a multi-institutional, multi-stage user study focusing on detecting mitoses in tumor images. Our analysis revealed that even with groups of three doctors, the majority-voting decisions had a higher appropriateness of AI reliance, compared to one doctor collaborating with AI. Subsequently, the majority voting decisions demonstrated increased precision and recall, although no statistical significance in recall was observed. Additionally, majority voting decisions were more likely to achieve complementary team performance in the recall. While effective on its own, majority voting can also be used together with other techniques to enable appropriate AI reliance. Involving multiple experts in decision-making can yield higher-quality, more robust outcomes that are less prone to AI errors, which holds promise in pathology and broader high-stakes domains.

### v3

@@ 4.5.1 Utilization of AI & XAI (RQ1)
    XAI activation rate was calculated according to Equation 3. The number of “AI recommendations in imagei” was counted based on the highest sensitivity set by a participant while they examined the imagei. If the “Show AI” was not toggled on in an image, then it was not counted.
@@ 4.5.2 Reliance on AI (RQ2)
Items Majority Voting Decisions (Group Size=$k$)
Human Decision (stage 1) Majority voting results based on the stage 1 decisions from $k$ participants
AI Recommendation (stage 2) For each image, AI recommendations under the highest sensitivity set by more than $k/2$ of participants while they were seeing the ROI
Human-AI Decision (stage 2) Majority voting results based on the stage 2 decisions from the same $k$ participants
@@ 4.5.3 Correctness of Mitosis Detection (RQ3)
Here, we compare the precision and recall of five conditions: one-human-AI collaboration, and majority voting decisions from AI-assisted pathologists (group sizes $k$=3,5,7,9). Results of larger group sizes are reported in the supplemental material. Similar to the comparisons in the AI reliance metrics, for each of precision and recall, a Kruskal–Wallis and a follow-up post-hoc Dunn’s test with Bonferroni correction was employed to test the significance among the condition pairs.
Because our previous work showed AI achieved higher overall performance than all participants in stage 1 (Gu et al., 2024), the “complementary team performance” in this work refers explicitly to cases where the human+AI approach outperforms AI (RQ3, i.e., super-AI performance). Here, the AI operating point was selected based on the best threshold in the model validation process. For precision and recall, we defined the “success rate of achieving super-AI performance” using Equation 9. This equation was applied to both the one-human-AI collaboration, and majority voting decision conditions with group sizes $k$ ranging from 3 to 27.
@@ 5.1 Utilization of AI & XAI
The mean AI activation rate was $M=99.21\%$ ($SD=0.481\%$, $CI_{95}=[98.13\%,100.00\%]$, Figure 6(a)) [^6] The mean ($M$), standard deviation ($SD$), and 95% confidence intervals ($CI_{95}$) were calculated by the bootstrapping method (100% re-sampling with replacement, 10,000 times) . And the mean AI active time percentage was $M=71.39\%$ ($SD=6.713\%$, $CI_{95}=[57.75\%,83.94\%]$, Figure 6(b)). 21/29 participants had $>50\%$ AI active time percentages, with an example of how they interacted with the “Show AI” feature shown in Figure 6(c), which suggests the user kept the AI activated for the majority of the time, with occasional brief flickering between turning it off and on during the initial interactions. The remaining 8/29 participants had $<25\%$ AI active time percentages: Although the “Show AI” feature was majority deactivated, these participants would still activate AI recommendations briefly while examining each image (Figure 6(d)). Interestingly, this pattern matches the cognitive forcing function (Buçinca et al., 2021) although these participants had not been instructed to do so.
@@ 5.3 Correctness of Mitosis Detection
Figure 8: Box-whisker plots of precision and recall for the five conditions of one-human-AI collaboration, and majority voting decisions ($k=3,5,7,9$); (c) Precision-recall plots for mitosis detection for these five conditions. The red line represents the precision-recall curve of AI, and the ‘x’ marker indicates the AI’s performance at a threshold determined by the best validation performance. The success rates of achieving super-AI
performance (i.e., percentage of human+AI cases where their performance is higher than both humans and AI) for the criteria of (d) precision and (e) recall.
@@
Regarding the success rates for achieving super-AI performance, for precision, none of the majority voting conditions (i.e., $k=3\rightarrow 27$) was higher than the success rate achieved by one-human-AI collaboration ($13.87\%$ success rate, Figure 8(d)). On the other hand, for recall, all majority voting conditions had higher success rates compared to one-human-AI collaboration ($51.72\%$ success rate): As shown in Figure 8(e), the lowest success rate was observed at $k=3$ ($53\%$ success rate), and the highest was achieved at $k=27$, reaching $76\%$.
@@ 6.1.1 Summary of RQ1
For most participants, AI was activated at least once in most images. However, this does not imply that the AI was constantly active throughout the entire study. Notably, 8/29 participants deactivated AI for most of the study, and only activated it briefly occasionally. That is, in certain instances, the ‘AI on-request’ feature posed cognitive forcing function effects.
@@ 6.1.3 Summary of RQ3
All majority voting conditions for $k=3\rightarrow 27$ did not show a higher success rate in achieving super-AI precision than one-human-AI collaboration. This is because AI had a high precision of 0.961, and there was a lack of space for improvement. For recall, all majority voting conditions ($k=3\rightarrow 27$) showed higher success rates. Notably, the highest success rate, 76%, was achieved at $k=27$, indicating a 46.95% increase over the one-human-AI collaboration condition (51.72% success rate).
Figure 9: Bar plots for the agreements among 29 participants for (a) 88 ground-truth mitoses, and (b) 91 false-positive mitoses that at least two participants agreed on. The diamond markers ($\lozenge$) stand for the AI detections under the “Highest” AI sensitivity setting. ① An example of under-reliance that might not be addressed by the majority voting; ② An example of under-reliance that might be addressed by the majority voting; ③ An example of over-reliance that might not be addressed by the majority voting; and ④ An example of over-reliance that might be addressed by the majority voting.
@@ 6.2 The Mechanism and Cost of Majority Voting
    Reducing Over-Reliance on AI False Positives: AI’s false-positive detections led to higher agreement rates among participants (as shown in Figure 9(b)③), suggesting participants’ tendency of over-reliance in at stage 2. The majority of these false-positive detections did not achieve agreement rates higher than 50% (Figure 9(b)④). In other words, from a group’s perspective, it was not usual for the majority of participants to consistently over-rely when AI made false-positive mistakes. Therefore, the over-reliance can be reduced by the majority voting mechanism.
@@
    Reducing Under-Reliance on Human False Positives: At stage 2, participants may make the same false-positive mistake as in stage 1, even when AI correctly suggested negative (Figure 9(b)①). This suggests that the under-reliance incidents happened when one participant collaborated with AI. Nevertheless, agreement rates for these false positives rarely exceeded the 50% majority threshold (Figure 9(b)②), indicating that majority voting could reduce under-reliance.
@@ 6.4 Towards Efficient & Reliable Medical Decisions with AI+k
Determining the optimal balance between decision-making and time expenditure has been well-explored in previous crowd-sourcing works (Daniel et al., 2018). However, one should be aware that the workflow of medical professionals is usually different from that of general users, and their preferences in using AI and XAI may also vary (as shown in Section 5.1). Therefore, future research should focus on exploring which AI+$k$ methods can seamlessly integrate into the workflow of medical professionals, effectively balancing efficiency and reliability in the medical decisions of multiple doctors. Additionally, investigating the role of counterfactual explanations (Zhou et al., 2022; Del Ser et al., 2024) to build trust and facilitate appropriate AI reliance could complement approaches like majority voting, potentially improving interpretability and familiarity with the decision process when integrating AI into risk-sensitive medical workflows.
@@ 7 Conclusion
This study introduces and validates the majority voting approach to enable doctors’ appropriate reliance on medical AI. By recruiting 32 pathology professionals, we conducted a multi-institutional, multi-stage user study focusing on detecting mitoses in tumor images. Our analysis revealed that even with groups of three doctors, the majority-voting decisions had a higher appropriateness of AI reliance, compared to one doctor collaborating with AI. Subsequently, the majority voting decisions demonstrated increased precision and recall, although no statistical significance in recall was observed. Additionally, majority voting decisions were more likely to achieve super-AI performance in the recall. While effective on its own, majority voting can also be used together with other techniques to enable appropriate AI reliance. Involving multiple experts in decision-making can yield higher-quality, more robust outcomes that are less prone to AI errors, which holds promise in pathology and broader high-stakes domains.